1. (8 pt) Solve the following ordinary differential equations by separate variables: dy 3x²+1 dx XyVit v2: y(1) = 0
Answer #1

7:14 PM 90 26%D < + $0 Ford xy vity? TOP0+ dy = 3x*! ,000=0 Separate the variables Sy ritge dy = (32+1 de > foregte poly = f*